OpenStudy (serenity2001):

Solve each equation giving a general formula for all of the solutions. 1. 2sin^2 x - 5sinx - 3 = 0 2. (cosx - √(2)/2)(secx - 1) = 0 I think the answer for #1 is x = 7π/6 + 2πn, x = 11π/6 + 2πn #2 are x = 7π/4 + 2πn

OpenStudy (daniel.ohearn1):

How did you factor it?

OpenStudy (daniel.ohearn1):

Only one answer for 1.

OpenStudy (daniel.ohearn1):

two answers for 2.

OpenStudy (daniel.ohearn1):

You can substitute sin(x) for a variable and factor the resulting equation right?

OpenStudy (serenity2001):

\[2\sin ^{2}x-5\sin x-3=0\] \[2u ^{2}-5u-3=0\] Like this, and then solve it?

OpenStudy (daniel.ohearn1):

yes

OpenStudy (daniel.ohearn1):

The 2nd one is already factored so solve each factor for zero using inverse functions.

OpenStudy (serenity2001):

I still got x = 7π/6 + 2πn and x = 11π/6 + 2πn for #1

OpenStudy (daniel.ohearn1):

Right, there's one factor two answers involving the sine function. Did you get the second one?

OpenStudy (daniel.ohearn1):

one useful factor for the first but u got it.

OpenStudy (serenity2001):

I got x = 7π/4 + 2πn and x = π/4 + 2πn